OpenStudy (anonymous):

Prove that log4(9) is irrational. I have no idea where to even begin on this.

OpenStudy (saifoo.khan):

Do you have a calc?

OpenStudy (anonymous):

Yeah, but sadly it's log function is all log10

OpenStudy (saifoo.khan):

No worries, use this rule.\[\Large \log_a b = \frac{\log b}{\log a}\]

OpenStudy (anonymous):

Really! That's impressive.

OpenStudy (saifoo.khan):

;)

OpenStudy (anonymous):

That should do me quite fine, I think I've got it from here.

OpenStudy (saifoo.khan):

Cool. B-)